OpenStudy (crashonce):

If f(2x) - 16x^3 + 20x, then what is f(3x)?

OpenStudy (anonymous):

do u mean f(2x) = 16x^3 + 20x

OpenStudy (crashonce):

ye

OpenStudy (anonymous):

let v=2x x=v/2 then f(v) - 16(v/2)^3 + 20(v/2)

OpenStudy (anonymous):

typo f(v) = 16(v/2)^3 + 20(v/2)

OpenStudy (anonymous):

got the idea ?

OpenStudy (anonymous):

so you would have this function f(v) = 16(v/2)^3 + 20(v/2) find when v=3x :)

hartnn (hartnn):

1. from f(2x), replace x by x/2 to get f(x) 2. from f(x), replace x by 3x to get f(3x) or you could combine both the steps above to : replace x by 3x/2 to get f(3x) from f(2x)

OpenStudy (crashonce):

@hartnn how would it go, can u show me

hartnn (hartnn):

go with easier method first... find f(x) from f(2x) by replacing all 'x' in f(2x) by 'x/2'

hartnn (hartnn):

like if i say replace x by \(\heartsuit\) in f(2x) you would do : \(f (2\heartsuit) = 16\heartsuit^3 + 20 \heartsuit \)

hartnn (hartnn):

similarly, replace x by x/2 instead

OpenStudy (crashonce):

54x^3+30x is final

hartnn (hartnn):

\(\huge \checkmark \)

hartnn (hartnn):

good work!